A newly designed wind turbine has three, 75-m long blades. On a sunny day, the turbine is operating such that it is rotating at 11 rev/min. If the tips of the blades follow a circular path, what is the speed of the tips of the blades as the turbine rotates?
Answer #1

Angular speed is

= 11 rev/min

= 11(2π rad)/(60 s)

= 1.152 rad/s

Speed of tip of blade is given by

= 75 × 1.152 = 86.4 m/s

(As tip rotates in a circle of radius 75 m)